a student throws a rock upwards. The rock reaches a maximum height 2.4 seconds after it was released. How fast was the rock thrown?
1
Expert's answer
2020-10-30T13:18:15-0400

We can find the initial velocity of the rock from the kinematic equation:


"v=v_0+gt,"

here, "v=0\ \dfrac{m}{s}" is the final velocity of the rock at the maximum height, "v_0" is the initial velocity of the rock, "g=-9.8\ \dfrac{m}{s^2}" is acceleration due to gravity and "t=2.4\ s" is the time that the rock takes to reach the maximum height.

Then, from this equation we can calculate the initial velocity of the rock:


"v_0=-gt=-(-9.8\ \dfrac{m}{s^2})\cdot 2.4\ s=23.52\ \dfrac{m}{s}."

Answer:

"v_0=23.52\ \dfrac{m}{s}."